[Brahms-l] Shift report 20060616 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 16 2006 - 07:55:26 EDT
Shifter leader was CN
Shifters were 


1. Had beam for 10 min :(. Lots of problems not allowing the MCR to keep the beam.

2. Took pedestals, and a single 10 min. data run

3. None

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 16 07:55:56 2006

This archive was generated by hypermail 2.1.8 : Fri Jun 16 2006 - 07:56:05 EDT